(3x^2y^3e^x+y^3+y^2)DX+(x^2y^3e^x-xy)sy=0


1
Expert's answer
2021-03-12T01:29:55-0500

"(3x^2y^3e^x+y^3+y^2)dx+(x^2y^3e^x-xy)dy=0\\
M(x,y)=3x^2y^3e^x+y^3+y^2\\
N(x,y)=x^2y^3e^x-xy\\
\frac{\partial M}{\partial y}=9x^2y^2e^x+3y^2+2y\\
\frac{\partial N}{\partial x}=2xe^xy^3-y\\
\frac{\partial M}{\partial y}\neq \frac{\partial N}{\partial x}\\
\frac{\frac{\partial M}{\partial y}-\frac{\partial N}{\partial x}}{N-M}=\frac{9x^2y^2e^x+3y^2+3y-2xe^xy^3}{-2x^2y^3e^x-xy-y^3-y^2}\\"